OpenStudy (anonymous):

Help..Find the center and radius: 2x^2 + 2y^2 + 8x + 7 = 0

OpenStudy (anonymous):

center (4, 0 )

OpenStudy (anonymous):

radius is 3

OpenStudy (anonymous):

are you sure ? maybe im too tired but i get this : x^2 + y^2 + 4x = -3.5 x^2 + 4x + 4 - 4 + y^2 = -3.5 (x+2)^2 + y^2 = -3.5 + 4 (x+2)^2 + y^2 = 0.5 center -2,0 radius (0.5)^0.5 = 1/(2^0.5)

OpenStudy (anonymous):

hmm are you sure ?

OpenStudy (anonymous):

my bad i didn't realize co efficient of x and y are also there

OpenStudy (anonymous):

i might be wrong but i dont see any mistake

OpenStudy (anonymous):

no you are right my mistake i should have divided it by 2 first

OpenStudy (anonymous):

sorry precalc

hero (hero):

Wow

OpenStudy (anonymous):

coolsector is right

OpenStudy (anonymous):

thank you both :)

hero (hero):

Well, at least it's solved

hero (hero):

Cool Sector, very good work